Sue had a total of $2500.00 to invest, She put part of it in an investment that pays 6% a year. The remaining money she puts in an investment that pays 5% a year. At the end of the year she has earned $138.50 in interest. How much money did she put in each of the investments.
Answer by mananth(16949)

Sue had a total of $2500.00 to invest, She put part of it in an investment that pays 6% a year. The remaining money she puts in an investment that pays 5% a year. At the end of the year she has earned $138.50 in interest. How much money did she put in each of the investments.
...
6%..................... $x
5% .....................$ 2500-x
earned $138.50 in interest.
Sue had a total of $2500.00

6x+5(2500-x)=138*100
6x+12500-5x=13800
x=13800-12500
x=1300
$1300 at 6%
$1200 at 5%
...
m.ananth@hotmail.ca